Please help me solve this problem!

Suppose a single firm produces all of the output in a contestable market. The market inverse demand function is P = 200 -2Q, and the firm’s cost function is C(Q) = 8Q. Determine the firm’s equilibrium price and corresponding profits.

Price: $

Profits: $
Answer #1

According given question the firm is monopolist

At maximum price

MR=MC

Given

P=200-2Q ;cost function C(Q)=8Q

TR=P×Q=(200-2Q)Q=200Q-2Q^2

MR=dTR/dQ=200-4Q

TC=8Q

MC=dTC/dQ=8

At maximizing MR=MC

200-4Q=8

4Q=200-8

   Q=192/4=48

Price=200-2Q=200-2×48=200-196=$104

profit=total revenue-total cost =104×48-8×48=4992-384=$4608